OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x – 6y = 12 9x + 2y = –24

OpenStudy (amistre64):

i wouldnt use the graphing method to solve any system of equations; id simply use it as a visual aid to get me in the ballpark

OpenStudy (anonymous):

Geometry without algebra is dumb (ask the Greeks), algebra without geometry is blind (ask the Chinese).

OpenStudy (anonymous):

could somebody just help me learn to graph it

OpenStudy (amistre64):

do you need all the points of each equation? or will just 2 for each one work?

OpenStudy (anonymous):

Ok, we just want to know whether u are expected to do it all by graph only.

OpenStudy (anonymous):

I have to do it all by graph. and two for each one will work!

OpenStudy (anonymous):

just explain how to graph it.... then help me solve the system....and explain how the solution represents the graph

OpenStudy (amistre64):

which 2 points do you think will be the easiest to find with these equations?

OpenStudy (amistre64):

2 points for each equation that is

OpenStudy (anonymous):

Do u know that they are straight lines?

OpenStudy (anonymous):

yes they are straight lines. @amistre64 and a point on the x axis and a point on the y axis?

OpenStudy (amistre64):

correct :) so we use x=0 in one case; and y=0 in the other and that should give us 2 points to draw a line with right?

OpenStudy (amistre64):

3x – 6y = 12 ; (x=0, y=?) ; (x=? , y=0)

OpenStudy (anonymous):

how would we solve the ? do you have to solve the 3x-6y=12 twice one time for y and another time for x?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

the equation is a road map; use it as many times as you need.

OpenStudy (anonymous):

(x=0, y=-2) ; (x=-3/-6 , y=0) is that it? im not quite sure

OpenStudy (amistre64):

if your not sure; how can we chk to make sure?

OpenStudy (anonymous):

plug it into the equation?

OpenStudy (amistre64):

thats right; your second point there is not going to fit into the equation very well, you might want to rethink it ..

OpenStudy (anonymous):

x=2y+4 so.... (x=0, y=-2) ; (x=4 , y=0)

OpenStudy (amistre64):

thats better :) now we can omit the "x=" and the "y=" parts to get our points looking like normal

OpenStudy (anonymous):

yes then we graph them....and are those points also the the solutions to the equation?

OpenStudy (amistre64):

those points are just solutions to ONE of the equations, namely the top equation.

OpenStudy (amistre64):

connect the points with a line that continues thru them

OpenStudy (anonymous):

oh okay. so 9x + 2y = –24; (x=0, y=12) ; (x=? , y=0). but when i solve for x 9 dosnt go into 24 so what would I do?

OpenStudy (amistre64):

your first point should be (0,-12) by the way fractions are fractions ... you best bet for an integer would be to solve for x in such a way that you can see an integer in it ... 9x + 2y = -24 ; -2y 9x = -24 - 2y ; /9 x = (-24 - 2y)/9 ;factor out a -1 x = -1 (24 + 2y)/9 now what numbers do 9 go into that we can make out of 24 + 2y ? 24 + 2(0) = 24 24 + 2(1) = 26 24 + 2(2) = 28 24 + 2(3) = 30 24 + 2(4) = 32 24 + 2(5) = 34 24 + 2(6) = 36 .. heres one ; 9(4) = 36 too so lets try this with y= 6 x = -1* (24 + 2(6))/9 = -1* 36/9 = -1* 4 = -4 ; (-4,6) looks like a good spot

OpenStudy (anonymous):

then for the solution i just find which point lies on both lines (intersection)